### v3

@@ Abstract
The polarization of light conveys crucial information about the spatial ordering and optical properties of a specimen. However, precise polarization measurement in challenging conditions, including constrained spaces, low light levels, and high-speed scenarios, remains a severe challenge. Addressing this problem, we introduce a real-time polarization measurement method that is accurate down to a single-photon level and provides complete information about the polarization state. Free of moving components, the polarization sensor utilizes a short rigid piece of few-mode fiber followed by a fiber array and a detector array. The calibration of the sensor relies on a neural network yielding unprecedented accuracy across all polarization states, including partially polarized light. We validate the approach by visualizing the polarization structure of biological specimens and the liquid crystal polymer sample (birefringent USAF test). Our method offers an efficient and reliable solution for real-time polarization sensing and microendoscopy under low-light conditions.
keywords:
light polarization, all-fiber sensor, in-situ and microendoscopic measurement, single-photon and low-light sensing, deep learning1 Introduction
@@
From imaging [1, 2] to sensing [3], from ellipsometry [4, 5] to optical communications [6, 7], light polarization represents a fundamental element across various scientific disciplines. In particular, the polarization of light emitted by an optically anisotropic specimen carries essential information about its material structure and optical properties. Consequently, conducting a thorough analysis of such specimens demands precise characterization of the emitted light polarization. While relatively straightforward in a controlled environment, accurate polarization measurement under challenging conditions remains a significant obstacle.
@@
Low light intensities, representing the first challenging condition, come with a reduced signal-to-noise ratio, which complicates extracting meaningful polarization information from data burdened with environmental and shot noise. Similarly, conducting high-speed measurements, as the second condition, is accompanied by low signal levels within the acquisition time. While increasing the incident light intensity can help, it is limited by our ability to deliver the optical power to the particular sensing area and, most ultimately, by the photodamage threshold of the sample. Incorporating single-photon detectors can substantially improve the accuracy of protocols relying on the number of detected photons for polarization state reconstruction. However, another drawback arises as these detectors are relatively large and cannot be efficiently integrated. This issue becomes particularly problematic when faced with the third condition - constrained space. Such spaces not only impose restrictions on the overall size of the measurement setup but also limit the placement and alignment of optical components. Therefore, these challenging conditions pose a severe setback, especially for measurements during invasive medical procedures, such as microendoscopy [8, 9], environmental monitoring [10, 11, 12], or internal/in-situ inspections of materials [13, 14]. We address these problems by developing a polarization sensor based on light propagation in disordered media.
@@
Coherent light propagating within a disordered medium is subjected to multimode interference, generating a granular speckle pattern. Despite the apparent degrading effect, the pattern retains encoded information about the incident light [15, 16, 17, 18, 19]. Analyzing the light propagation in disordered media and employing suitable post-processing techniques, one can study the disordering effects of the media [20, 21], estimate and classify input images [22, 23, 24, 25, 26, 27, 9], learn to modulate the input for projecting a targeted image on the output [28], reconstruct shapes of propagated ultrashort laser pulses [29, 30], compute and communicate [31, 32, 33, 34, 35], and even make an accurate spectrometer or wavemeter [36, 37]. Recently, the speckle pattern has proved to preserve a great deal of information about the polarization state of the incident light [38, 39]. Although additional studies successfully advancing the application of disordered-media propagation have been conducted [40, 41, 42, 43], developing a compact polarimeter capable of rapid measurements with single-photon sensitivity continues to be an open problem.
@@
Here, we introduce a highly accurate single-shot all-fiber polarization sensor based on intermodal interference in a very short piece of a few-mode fiber. Instead of processing the complete spatial speckle pattern, our approach relies on its sparse sampling in a few isolated points using a fiber array. Fiber-coupled single-photon detectors then allow for high sensitivity down to a single-photon level. The all-fiber sensor contains no moving components or complex metastructures and is sufficiently small for endoscopic procedures. Furthermore, the method can accurately operate at a high repetition rate, processing several thousands of polarization states per second. The setup is aided by a deep machine-learning model reconstructing complete polarization information in real time. The calibrated sensor maintains unparalleled accuracy and stability throughout an extensive period, even for partially-polarized light. We demonstrate the sensing abilities of our approach in polarized light scans of dense connective tissue, moving diatom, and the birefringent USAF test. Our polarization sensor opens new frontiers in precise polarization measurements, providing an indispensable tool for cutting-edge applications in biomedicine, material research, and beyond.
2 Results and Discussion
2.1 Experimental and data processing setup
@@
The speckle-based polarization sensor depicted in Fig. 1 employs a common few-mode fiber to characterize the input polarization state. The sensor collects light with a front face of a 5 cm long step-index SMF28 fiber with a numerical aperture of 0.14 and a core diameter of 8.2 m. While the motivation behind choosing this specific fiber type is to minimize the sensing area and keep the speckle structures large, we emphasize that this preference is optional. Virtually any few-mode or multimode fiber, including significantly shorter variants, can be used as a substitute. The main requirement for the fiber is to support several interfering modes necessary for creating a polarization-dependent speckle pattern at the output. Following a 5mm free-space propagation, we collect discrete samples of this interference pattern using a fiber array comprising seven gradient multimode fibers, each with a 62.5 m core diameter, forming a 375 m diameter honeycomb pattern. Each fiber then propagates the collected intensity signal to an independent single-photon avalanche diode, recording the number of detection events. After normalizing the measured counts, this setup associates the sampled polarization state with the corresponding relative frequencies of each detector, which we term a count distribution. For more experimental details, see Sec. 4.1.
@@
Using the sensor, we analyzed 30,000 polarization states, uniformly covering the entire surface of the Poincaré sphere [44]. Our state preparation involved a continuous laser beam with a central wavelength of 0.8 m attenuated to the single-photon level and propagated through a linear polarizer. Subsequently, a twisted nematic liquid crystal device [45], controlled by voltage signals, performed a targeted unitary transformation of the polarization state on this weak coherent light. The device operation underwent calibration through bidirectional modeling [46], ensuring fast and precise preparation of arbitrary polarization states. Its accuracy was verified using a reference polarimeter based on wave plates. The polarization state preparation using both methods reaches an average fidelity exceeding 0.999, however, the liquid crystal device operates by orders of magnitude faster. Subsequently, our sensor characterized the prepared polarization state, providing the corresponding detection counts. On average, we acquired approximately 150.000 photon detection events per detector within a 50 ms acquisition time window. Moreover, we expand our dataset to include partially polarized light. We can establish a count distribution for a mixture of two orthogonal states by weight-summing their respective distributions. This approach is virtually equivalent to directly measuring the corresponding mixed state. Employing this method, we assembled an extensive dataset of polarization states distributed uniformly throughout the entire volume of the Poincaré sphere.
We used the dataset to train a fully-connected deep neural network for reconstructing the polarization state given the associated count distribution, see Sec. 4.3 for details. The optimized and fully-trained network was evaluated on a test set of previously unobserved polarization states, also acquired from the experimental setup. We achieved the unprecedented average infidelity of with a confidence interval using the 10th and 90th percentile. These results reflect the highest achievable measurement accuracy as they approach the fidelity of our polarization state preparation. Furthermore, we conducted additional validation of our results using a 1951 USAF test, as detailed in Sec. 4.4. This supplemental analysis provides further insight into the overall performance of our approach.
2.2 Performance evaluation
@@
We highlight that our approach eliminates the need for capturing an image of the complete speckle pattern. Instead, we collect a limited number of isolated samples, which are then propagated through the fiber array channels. This substantial reduction in data collection allows for sensitive measurements at extremely low intensities using single-photon detectors. To assess the sensitivity of our approach to the processed portion of the speckle pattern, we investigate the achievable infidelity with regard to the number of employed detection channels. Following the same procedure, we trained additional networks utilizing only a subset of the detection channels as inputs. The infidelities achieved by these networks on corresponding test sets are depicted in Fig. 2 (a). The values represent the averaged errors across multiple trained networks, each employing a distinct combination of detection channel subsets. The results illustrated in this graph unequivocally demonstrate that extracting only a few samples from the speckle pattern provides sufficient information about the polarization state.
@@
We also explored the dependence of the polarization sensing accuracy on the measurement repetition rate, i.e., the inverse value of the acquisition time, representing the number of polarization states characterized per second. Fig. 2 (b) displays the polarization infidelities of a single network, using all seven channels, evaluated for numerous repetition rates. Notably, achieving an average fidelity of 0.999 at a 33 Hz rate holds particular significance. This frequency aligns closely with the borderline on a perceptible frame rate of the human eye. Hence, it underscores the ability of the sensor to operate in real-time sensing applications with high accuracy. Furthermore, our setup demonstrates the capacity to measure over 2,000 polarization states per second while maintaining an average fidelity surpassing 0.99. It is crucial to emphasize that these performances depend on the number of detected photons rather than the repetition rate alone. Consequently, even significantly higher repetition rates can attain the same fidelities with appropriately increased light intensity limited virtually only by saturation of employed single-photon detectors. This feature proves especially beneficial for applications requiring high-repetition-rate sensing.
Furthermore, let us discuss the long-term stability of our approach and its resistance to environmental effects. Despite using a 5 cm long few-mode fiber in an open environment without elaborated fixation, the polarization sensor shows sufficient stability for a week without additional calibration or network retraining. For in-situ and microendoscopic applications, the fiber is typically inserted in a cannula, preventing bending and micro deformations. The stability can be further enhanced using a few-mm or sub-mm long fiber encased within a ceramic or glass ferrule tip. The millimeter-long fiber provides sufficient intermodal coupling, and the encasing fixes the fiber and its position relative to the fiber array. With these modifications, we expect an improvement in stability by up to two orders of magnitude. Employing very short few-mode fiber represents the major difference compared to imaging applications, where long and massively multimode fibers have to be used with deteriorating impact on the stability [22, 23, 24, 25, 26, 27, 9]. In summary, our polarization sensor represents an outstanding instrument for efficient and accurate real-time and high-speed polarimetry applications.
2.3 Applications
@@
We first demonstrate the sensing ability of our approach by conducting a polarization-sensitive scan of dense connective tissue (AmScope PS25W). While the primary application of our sensor significantly differs from polarization imaging, this optically anisotropic specimen is famously known in the polarization microscopy field. Visualizing its typical structures further underscores the accuracy of the presented polarimetric sensor. Using the experimental arrangement outlined in Sec. 4.1, we replaced the calibration polarization preparation stage with the tissue specimen. The few-mode fiber positioned proximal to the tissue scanned over its area, capturing the induced polarization transformation. The reconstructed polarization states were expressed as three Bloch parameters, i.e., renormalized Stokes parameters [44], and visualized as a false-colored RGB image, see Sec. 4.4 for details. In Fig. 3, we present a side-by-side comparison of this polarization scan (middle) with an intensity image (left) captured from an identical tissue region. This intensity image was obtained using a monochrome camera (ImagingSource DMK 23U274) with a pixel size of m. We illuminated the specimen with 0.8 m light and projected the resulting image onto the camera through a 25.4 mm focal-length lens, providing an estimated fivefold magnification. Moreover, the third panel (right) depicts a polarization image of the same specimen obtained using a stand-alone polarization microscope. Its experimental setup comprises a pair of crossed polarizers and a monochrome camera (ImagingSource DMK 23U274). As illustrated in Fig. 3, both polarization-based methods reveal underlying polarization structures that remain concealed in the intensity profile. This typical inner structure of the dense connective tissue verifies that our sensor performs polarization characterization, providing complete information about the polarization state.
@@
As the dense connective tissue is a static specimen, the complete polarization analysis provided by the all-fiber sensor can also be performed using an advanced imaging polarimeter comprising rotating wave plates. However, such a technique is insufficient for characterizing a rapidly changing polarization state induced by moving or evolving specimens. These scenarios require significantly faster approaches that can be performed in a single-shot regime. We demonstrate the feasibility of our sensor for such applications by conducting a polarization measurement to characterize the floating diatom. These microscopic unicellular organisms comprise a complex inner structure that can exhibit anisotropic properties. By continuously measuring the polarization state, we characterized the polarization transition caused by a moving diatom within the same experimental arrangement as in the previous application. The evolution of the Bloch parameters corresponding to this dynamic process is depicted in Fig. 4 with red lines. Compared to the blue lines of background polarization, the diatom properties are visible in the modulation of Bloch parameters. One could count the number of living specimens, estimate their size and shape, or even classify the detected organisms based on their birefringent properties. The sensor can also find applications for fast in-situ material inspection, local strain analysis, and micro crystal growth monitoring. These showcases underscore the versatility of our polarization sensor for various sensing applications and its valuable contributions to low-light high-speed polarimetry.
3 Conclusion

### v4

@@ Abstract
Polarization of light carries vital information in numerous scientific disciplines, including biomedical imaging, optical diagnostics, and environmental sensing. However, accurate polarization measurements in constrained spaces, under low-light conditions, and at high speeds remain a severe challenge. We present a compact, all-fiber polarization sensor capable of single-shot, real-time operation with single-photon sensitivity and long-term stability. The sensor leverages intermodal interference within a short segment of a few-mode fiber, coupled to an array of single-photon detectors. The recorded detections are processed by a neural network, enabling precise reconstruction of complete polarization information for fully and partially polarized states. This robust architecture allows for thousands of polarization state measurements per second while achieving exceptional accuracy with Stokes error below 0.01, and even lower at higher photon fluxes. We demonstrate this technology through diverse experimental scenarios, such as resolving structural details in biological tissues with a spatial resolution of 6 m, characterizing rapid polarization transitions, and monitoring micro-scale birefringence in living or moving specimens.
I Introduction
@@
From imaging [1, 2] and sensing [3] to ellipsometry [4, 5] and optical communications [6, 7], light polarization plays a fundamental role across numerous scientific disciplines. In particular, the polarization of light emitted by an optically anisotropic specimen carries essential information about its material structure and optical properties. This capability has led to widespread applications in medicine, where polarization provides insight into the structural and morphological characteristics of various biological tissues and enables label-free detection of pathological changes [8]. Polarization-resolved endoscopy and imaging techniques have proven valuable for detecting, classifying, and staging cancerous lesions in organs such as the stomach, colon, prostate, and larynx [9, 10]. However, while polarization measurements can be relatively straightforward in a controlled laboratory environment, achieving highly accurate and reliable characterization under challenging conditions and in complex dynamic scenarios remains a significant obstacle.
@@
Low light intensities, representing the first challenging condition under consideration, come with a reduced signal-to-noise ratio, which complicates extracting meaningful polarization information from data burdened with environmental and shot noise. Similarly, conducting high-speed measurements, as the second condition, is accompanied by low signal levels within the acquisition time. While increasing the incident light intensity can help, it is limited by our ability to deliver the optical power to the particular sensing area and, most ultimately, by the photodamage threshold of the sample. Incorporating single-photon detectors can substantially improve the accuracy of protocols relying on the number of detected photons for polarization state reconstruction. However, another drawback arises as these detectors are relatively large and cannot be efficiently integrated. This issue becomes particularly problematic when faced with the third condition - constrained space. Such spaces not only impose restrictions on the overall size of the measurement setup but also limit the placement and alignment of optical components. Therefore, these challenging conditions pose a severe setback, especially for measurements during invasive medical procedures, such as microendoscopy [11, 12], environmental monitoring [13, 14], or internal/in-situ inspections of materials [15, 16]. We address these problems by developing a polarization sensor based on light propagation in disordered media.
@@
Coherent light propagating within a disordered medium is subjected to multimode interference, generating a granular speckle pattern. Despite the apparent degrading effect, the pattern retains encoded information about the incident light [17, 18, 19, 20, 21]. Analyzing the light propagation in disordered media and employing suitable post-processing techniques, one can study the disordering effects of the media [22, 23], estimate and classify input images [24, 25, 26, 27, 28, 12], learn to modulate the input for projecting a targeted image on the output [29], reconstruct shapes of propagated ultrashort laser pulses [30, 31], compute and communicate [32, 33, 34, 35, 36], and even make an accurate spectrometer or wavemeter [37, 38]. Recently, the speckle pattern has proved to preserve a great deal of information about the polarization state of the incident light [39, 40]. Although additional studies successfully advancing the application of disordered-media propagation have been conducted [41, 42, 43, 44], developing a compact polarimeter capable of rapid measurements with single-photon sensitivity continues to be an open problem.
@@
Here, we introduce a highly accurate, single-shot, all-fiber polarization sensor based on intermodal interference in a very short segment of a few-mode fiber. Instead of processing the complete spatial speckle pattern, our approach relies on its sparse sampling in a few isolated points using a fiber array. Combined with fiber-coupled single-photon detectors, this design allows for high sensitivity down to a single-photon level. The all-fiber sensor contains no moving components or complex metastructures and is sufficiently compact for procedures in constrained spaces while operating at a spatial resolution of 6 m. Additionally, it supports a high-throughput operation capable of resolving several thousand polarization states per second. The setup is aided by a deep machine-learning model reconstructing complete polarization information in real-time. Once calibrated, the sensor maintains unparalleled accuracy, reaching Stokes errors well below 0.01, and long-term stability for over a month, even when measuring partially polarized light. We demonstrate its polarimetric capabilities through various experiments, including scans of dense connective tissue, birefringence measurements with a USAF test target, and dynamic polarization sensing in a moving diatom and during a fast twisted nematic liquid crystal transition. This all-fiber sensor sets new standards for compact, stable, and high-speed polarization measurements, offering a powerful tool for cutting-edge applications in biomedicine, material research, and beyond.
II Results
II.1 Experimental and data processing setup
@@
The speckle-based polarization sensor illustrated in Fig. 1 employs a common few-mode fiber to characterize the input polarization state. The sensor collects light with a front face of a 5 cm long step-index SMF28 fiber with a numerical aperture of 0.14 and a core diameter of 8.2 m. While the motivation behind choosing this specific fiber type is to minimize the sensing area and keep the speckle structures large, we emphasize that this preference is optional. Virtually any few-mode or multimode fiber, including significantly shorter variants (see Sec. II.2), can be used as a substitute. The main requirement for the fiber is to support several interfering modes necessary for creating a polarization-dependent speckle pattern at the output. Following a 5mm free-space propagation, we collect discrete samples of this interference pattern using a fiber array comprising seven gradient multimode fibers, each with a 62.5 m core diameter, forming a 375 m diameter honeycomb pattern. Each fiber then propagates the collected intensity signal to an independent single-photon avalanche diode, recording the number of detection events. After normalizing the measured counts, this setup associates the sampled polarization state with the corresponding relative frequencies of each detector, which we term a count distribution. For more experimental details, see Sec. IV.1 and Fig. 7.
@@
Using the sensor, we analyzed 30,000 polarization states, uniformly covering the entire surface of the Poincaré sphere [45]. Our state preparation involved a continuous laser beam with a central wavelength of 0.8 m attenuated to the single-photon level and propagated through a linear polarizer. Subsequently, a twisted nematic liquid crystal device [46], controlled by voltage signals, performed a targeted unitary transformation of the polarization state on this weak coherent light. The device operation underwent calibration through bidirectional modeling [47], ensuring fast and precise preparation of arbitrary polarization states. Its accuracy was verified using a reference polarimeter based on wave plates. The polarization state preparation using both methods reaches an average fidelity exceeding 0.999, however, the liquid crystal device operates by orders of magnitude faster. Subsequently, our sensor characterized the prepared polarization state, providing the corresponding detection counts. On average, we acquired approximately 150.000 photon detection events per detector within a 50 ms acquisition time window. Moreover, we expand our dataset to include partially polarized light. We establish a count distribution for a mixture of two orthogonal states by weight-summing their respective distributions. This approach is virtually equivalent to directly measuring the corresponding mixed state. Employing this method, we assembled an extensive dataset of polarization states distributed uniformly throughout the entire volume of the Poincaré sphere.
We used the dataset to train a fully connected deep neural network for reconstructing the polarization state given the associated count distribution, see Sec. IV.3 for details. The optimized and fully-trained network was evaluated on a test set of previously unobserved polarization states, also acquired from the experimental setup. We achieved an unprecedented average infidelity of with a confidence interval using the 10th and 90th percentile. These results closely approach the fidelity limit imposed by the polarization preparation itself [46], indicating that our method reaches the fundamental accuracy bounds of the experiment. The achieved infidelity translates to a reconstruction error of approximately 0.01 per Stokes parameter. In contrast, previous methods based on random scattering typically reach errors around 0.05 while requiring bulky setups unsuitable for in-situ applications, achieving lower operational speed, lacking the single-photon sensitivity, and not presenting the ability to process partially polarized light [39, 40, 41, 42]. To further validate our approach, we performed an additional benchmark using a 1951 USAF birefringent test. As detailed in Sec. IV.4, we compared our sensor with a polarization measurement based on rotating wave plates. The two methods yielded an average fidelity of approximately 0.987, underscoring the accuracy of our all-fiber sensor. Additionally, the USAF test also allowed the sensor spatial resolution estimation, which was determined to be approximately m. Together, these results comprehensively demonstrate the exceptional accuracy and robustness of our approach.
II.2 Performance evaluation
@@
We highlight that our approach eliminates the need for capturing an image of the complete speckle pattern. Instead, we collect a limited number of speckle spots, which are then propagated through the fiber array channels. This substantial reduction in data collection allows for sensitive measurements at extremely low intensities using single-photon detectors. To assess the sensitivity of our approach to the processed portion of the speckle pattern, we investigate the achievable infidelity with regard to the number of employed detection channels. Following the same procedure, we trained additional networks utilizing only a subset of the detection channels as inputs. The infidelities achieved by these networks on corresponding test sets are depicted in Fig. 2 (a). The values represent the averaged errors across multiple trained networks, each employing a distinct combination of detection channel subsets. The results illustrated in this graph unequivocally demonstrate that extracting only a few samples from the speckle pattern provides sufficient information about the polarization state.
@@
We also explored the dependence of the polarization sensing accuracy on the measurement repetition rate, i.e., the inverse value of the acquisition time, representing the number of polarization states characterized per second. Fig. 2 (b) displays the polarization infidelities of a single network, using all seven channels, evaluated for numerous repetition rates. This analysis allows us to determine the minimum required light intensity (i.e., the number of detected photons) needed to achieve a targeted infidelity value. Similarly, it allows estimation of the attainable fidelity for a given repetition rate. For example, the system achieves an average fidelity of 0.999 at a rate of 33 Hz - approximately the threshold at which the human eye perceives continuous motion. This result highlights the capability of our sensor for accurate, real-time polarization measurements. Furthermore, our setup demonstrates the capacity to measure over 2,000 polarization states per second using only photons while maintaining an average fidelity above 0.99. It is crucial to emphasize that these performances depend on the number of detected photons rather than the repetition rate alone. Consequently, even significantly higher repetition rates can attain the same fidelities with appropriately increased light intensity, limited virtually only by saturation of employed single-photon detectors. The presented photon efficiency makes the sensor particularly suitable for high-throughput, rapid-sensing applications.
Furthermore, let us discuss the long-term stability of our approach and its robustness against environmental effects. Even when using a 5 cm long, few-mode fiber in an open environment without elaborated mechanical stabilization, the polarization sensor maintained sufficient stability for a week without recalibration or network retraining. We explored the possibility of further enhancing the stability using a 12 mm long fiber encased within a ceramic ferrule. This encasing rigidly fixes the fiber position relative to the fiber array, minimizing mechanical bending and reducing susceptibility to temperature gradients. With this modification, the system maintained its performance for over a month without recalibration. Fig. 3 illustrates the results of our long-term stability test, which was evaluated periodically at 24-hour intervals. The average fidelity was observed to degrade by approximately per day, i.e., per hour. This exceptional stability is attributed to three key factors: the short fiber length, which minimizes propagation-induced instabilities; the ferrule encasing, which mitigates environmental perturbations; and the use of a few-mode fiber, which produces a coarse-grained speckle pattern that increases the collection efficiency and reduces sensitivity to vibrations and drift. This design starkly contrasts with imaging-based applications, where long and massively multimode fibers have to be used with a deteriorating impact on the stability [24, 25, 26, 27, 28, 12]. Since sufficient intermodal coupling to generate a speckle pattern in a multimode fiber occurs over a propagation distance of several hundred wavelengths [48, 49], we could reduce the employed fiber length even to sub-millimeter lengths. We expect that reducing the fiber to this scale could yield an additional order-of-magnitude improvement in stability. Finally, the sensor can be recalibrated to prolong its stability indefinitely. The recalibration process is straightforward, fully automatable, and can be completed in under an hour, making regular, even daily, recalibration a practical option. In summary, our polarization sensor offers a highly stable, accurate, and efficient solution for real-time, high-speed polarimetry with minimal maintenance requirements and strong resistance to environmental disturbances.
II.3 Applications
@@
We first demonstrate the sensing ability of our approach by conducting a polarization-sensitive scan of dense connective tissue (AmScope PS25W). While the primary application of our sensor significantly differs from polarization imaging, this optically anisotropic specimen is famously known in the polarization microscopy field. Visualizing its typical structures further underscores the accuracy of the presented polarimetric sensor. Using the experimental arrangement outlined in Sec. IV.1, we replaced the calibration polarization preparation stage with the tissue specimen. The few-mode fiber was positioned proximal to the tissue, capturing the induced polarization transformation. We used motorized translation stages (Newport MFA-CC) with a bidirectional repeatability of m and a maximum speed of m/s to scan over the tissue area. The reconstructed polarization states were expressed as three Bloch parameters, i.e., renormalized Stokes parameters [45], and visualized as a false-colored RGB image, see Sec. IV.4 for details. In Fig. 4, we present a side-by-side comparison of this polarization scan (middle) with an intensity image (left) captured from an identical tissue region. This intensity image was obtained using a monochrome camera (ImagingSource DMK 23U274) with a pixel size of m. We illuminated the specimen with 0.8 m light and projected the resulting image onto the camera through a 25.4 mm focal-length lens, providing an estimated fivefold magnification. Moreover, the third panel (right) depicts a polarization image of the same specimen obtained using a stand-alone polarization microscope. Its experimental setup comprises a pair of crossed polarizers and the monochrome camera. As illustrated in Fig. 4, both polarization-based methods reveal underlying polarization structures that remain concealed in the intensity profile. This typical inner structure of the dense connective tissue verifies that our sensor performs polarization characterization, providing complete information about the polarization state.
@@
As the dense connective tissue is a static specimen, the complete polarization analysis provided by the all-fiber sensor can also be performed using an advanced imaging polarimeter comprising rotating wave plates. However, such a technique is insufficient for characterizing a rapidly changing polarization state induced by moving or evolving specimens. These scenarios require significantly faster approaches that have to be performed in a single-shot regime. To demonstrate the dynamic capabilities of our all-fiber sensor, we conducted polarization measurements of two systems: a floating diatom and twisted nematic liquid crystals undergoing a fast voltage-controlled transition. Starting with diatoms, these microscopic unicellular organisms comprise a complex inner structure that can exhibit anisotropic properties. We characterized the birefringence of an Actinoptychus heliopelta diatom positioned within the same experimental arrangement as the previous application. The spatial variation in the Bloch parameters across the diatom is depicted in Fig. 5 with red lines. Compared to the blue lines of background polarization, the diatom properties are visible in the modulation of Bloch parameters. This analysis can count the number of living specimens, estimate their size and morphology, and classify them based on their birefringent properties. Beyond biological samples, the sensor can also find applications for fast in-situ material inspection, local strain analysis, and micro crystal growth monitoring.
Finally, we demonstrate the high-speed polarimetric capabilities of our sensor using a rapidly evolving specimen: twisted nematic liquid crystals. They allow for fast polarization state transformation, controlled by low-voltage signals that manipulate their spatial alignment. During reorientation, the induced polarization transformation undergoes a rapid transition, typically within a time window ranging from 50 to 600 ms [46]. We characterized the horizontal-to-vertical transition by continuously measuring the prepared polarization state using a 5 ms acquisition window. The reconstructed Bloch parameters corresponding to this process are visualized in Fig. 6. The first Block parameter exhibits a smooth evolution from (horizontal polarization) to (vertical polarization). The remaining two parameters reveal that the transition corresponds to a continuous rotation on a Bloch sphere, passing close to the diagonal state (). The complete transition occurs within approximately 50 ms, validating the ability of our sensor to accurately capture dynamic polarization changes in real time, which is inaccessible to common rotating-wave-plate polarimeters. This demonstration, alongside the previously discussed applications, underscores the versatility and performance of our all-fiber polarization sensor for fast, low-light, high-precision polarimetry.
III Discussion